OpenStudy (anonymous):

Formulas for Q3? "For the last two parts assume I = 2A" Also, formulas for Q6 please :)

OpenStudy (anonymous):

q6 the formula 1-Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2-i1=V0/(R3-B*R1*R3) 3- vo = vth but i don't know the RTH

OpenStudy (anonymous):

Rth plzzzz

OpenStudy (anonymous):

Thank you Sara! I'm still missing Rth formula, and the 2nd part of Q3 ones...

OpenStudy (anonymous):

When I=2 v1 is 1,32 and v2 is 0,66

OpenStudy (anonymous):

Thank you zozi, but that seems to be wrong:c

OpenStudy (anonymous):

what formulas did you use? My Resistances are 1ohm

OpenStudy (anonymous):

just valculating v1, one 1 ohm and two serial resistors in paralel means a 0,66ohm resistor. From this U=I*R for v1 and for v0 its a half divider from two one ohm resistor...

OpenStudy (benzo):

sara do you mean I1= vin/(r3-b*r1*r3 vo=vin*(r3-(b*r1*r3)+r1+r2+a*r2)/(r3-b*r1*r3)

OpenStudy (benzo):

q6

OpenStudy (anonymous):

Q6 Rth please Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5.

OpenStudy (anonymous):

como vc fez o Q3, Joao?

OpenStudy (anonymous):

eu achei aki as respostas para o meu

OpenStudy (anonymous):

Thank you once again zozi, it was v0= 0.666 and v1=1.32, it was the other way around Also, for you Joao Vitor, RTH = 3.6

OpenStudy (benzo):

jones what the formula for q6

OpenStudy (anonymous):

are you sure abt that, Jone133?

OpenStudy (anonymous):

Yes, for I=2A on Q3 and for the values of JoaoMC on Q6

OpenStudy (anonymous):

THANK YOU SOO MUCH!!

OpenStudy (anonymous):

RTH= Vth / ISC Isc= Current on short circuit

OpenStudy (benzo):

I1 Vo Vth Rth

OpenStudy (anonymous):

but how did you calculate the Rth ???

OpenStudy (anonymous):

Finally 100% \o/

OpenStudy (anonymous):

congrats! It's a great accomplishment, I hope I get 100% too on the final :D

OpenStudy (benzo):

formula please jones...

OpenStudy (anonymous):

Hey Jone133.. my values are: 10A and -3A for v0 and 5A for v1 and v0

OpenStudy (anonymous):

Dá uma força com o 6 aí, joao?? =D

OpenStudy (anonymous):

I don't have formulas, sorry... I Only know that RTH = VTH / Isc and Isc is the current on short circuit. Zizo told me the values for Q3.

OpenStudy (benzo):

mc congrats q6 whats the formula

OpenStudy (anonymous):

para os valores de JoaoMC, Vo, I and Vth are 1.8 The formulas are on the 1st answer

OpenStudy (benzo):

ok I1 and vo

OpenStudy (anonymous):

also, vo = vth

OpenStudy (anonymous):

Q6: Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5. 1.8 1.8 1.8 3.6

OpenStudy (anonymous):

but still no formulas...

OpenStudy (anonymous):

For Q3 v0= 0,66*I/2 but if its over +/- 2V the zener cuts it to +/-2V (found zizo's answer on another question)

OpenStudy (anonymous):

Thank you Zozi!

OpenStudy (anonymous):

and v1 is 0,66*I

OpenStudy (benzo):

JoãoVitorMC u is the mannn thank u

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

i will post soon a topic here about solve all the Question 3 and others that i complete understant.

OpenStudy (anonymous):

Eu vi que teu Rth deu o dobro do teu Vth.. posso assumir isso pro meu, será? meu vth tá em 7.5

OpenStudy (anonymous):

acho melhor não arriscar não

OpenStudy (anonymous):

espera até amanha, deve aparecer mais soluções

OpenStudy (anonymous):

agora q acabei aki, vou trabalhar em como encontrar as respostas e irei postar aki o mais rapido possivel

OpenStudy (anonymous):

Q3: p1 : zener voltage sholud be less than the given current so vo=2V which is zener voltage Q3: P2 : I1=( 0.66 * I ) . . . also find I2 = (0.33 * I) .. . Vo= I2 Q3: P3 : (a) I1=( 0.66 * I ) . . . also find I2 = (0.33 * I) .. . Vo= I2 . . (b) v1 = I1

OpenStudy (anonymous):

what is Isc in Q6

OpenStudy (anonymous):

Question 3 explained: http://openstudy.com/study#/updates/5089d2f4e4b077c2ef2e01ca

OpenStudy (anonymous):

Rth help plzzz

OpenStudy (anonymous):

if you assume I=2A then the current through the parallel resistor will be 1.3333 A ant V1 will be(1ohm*1.333) 1.333V the remaining current flows through the zener diode its resistance will be 1ohm so v0 will be 0.667v

OpenStudy (anonymous):

anyone qno 4 part first equation????